# Latent Inversion with Timestep-aware Sampling for Training-free Non-rigid Editing

### v2

@@ Abstract
Text-guided non-rigid editing involves complex edits for input images, such as changing motion or compositions within their surroundings. Since it requires manipulating the input structure, existing methods often struggle with preserving object identity and background, particularly when combined with Stable Diffusion. In this work, we propose a training-free approach for non-rigid editing with Stable Diffusion, aimed at improving the identity preservation quality without compromising editability. Our approach comprises three stages: text optimization, latent inversion, and timestep-aware text injection sampling. Inspired by the recent success of Imagic, we employ their text optimization for smooth editing. Then, we introduce latent inversion to preserve the input image’s identity without additional model fine-tuning. To fully utilize the input reconstruction ability of latent inversion, we suggest timestep-aware text inject sampling. This effectively retains the structure of the input image by injecting the source text prompt in early sampling steps and then transitioning to the target prompt in subsequent sampling steps. This strategic approach seamlessly harmonizes with text optimization, facilitating complex non-rigid edits to the input without losing the original identity. We demonstrate the effectiveness of our method in terms of identity preservation, editability, and aesthetic quality through extensive experiments.
@@ 1 Introduction
Diffusion models have become a new standard in image generation and have had a significant impact on text-to-image generation tasks. Text-guided diffusion models take text prompts as input and generate images that align well with the input text prompt. Building upon the success of text-guided diffusion models, there has been active research in text-guided image editing. This encompasses global style transfer Tumanyan et al. (2023); Valevski et al. (2023); Wang et al. (2023b); Zhang et al. (2022) and local editing that require changing object attributes (e.g., color, shape), image translation and adding details Hertz et al. (2022); Couairon et al. (2022); Avrahami et al. (2022); Brack et al. (2023b).
Figure 1: Our method excels detailed non-rigid edits, such as changing the motion or composition of an object on real images directed by text prompts via Stable Diffusion. This figure showcases original images with their corresponding text-guided modification. Our training-free method preserves the original identity and context of the input images while achieving high-level conceptual alterations that seamlessly blend with the images’ inherent characteristics.
@@
Among various local editing tasks, non-rigid editing involves changing the posture or composition of an object while preserving the background and object identity of the original image. It is useful for various applications, such as drawing comic books or videos, where gradual transitions in object actions or compositions over the input image are essential. This task is particularly more challenging than rigid editing, such as global style transfer, image translation, and appearance transformation, which focus on changing color or textures while retaining the image structure. Instead, non-rigid editing requires modifying an object structure partially to align with the text prompt while maintaining its identity.
@@
The challenges of non-rigid editing become more severe when combined with Stable Diffusion Rombach et al. (2022), primarily related to preserving image identity. Prior works in image editing address this problem through various approaches, including injecting attention maps of the source image into the target image generation process Hertz et al. (2022); Tumanyan et al. (2023); Xiao et al. (2023); Cao et al. (2023), model finetuning Ruiz et al. (2023); Valevski et al. (2023); Kawar et al. (2023), and target text optimization Gal et al. (2022); Han et al. (2023); Kawar et al. (2023). However, the attention mechanism often leads to unnatural images as it forcefully combines the structure of the source image with the target, resulting in distorted objects and compositions. Also, model finetuning on an input image can lead to a severe overfitting problem and color distortion, especially when applied to Stable Diffusion.
@@
To this end, our work explores improving identity preservation for non-rigid editing with Stable Diffusion by incorporating a text optimization framework. We focus upon the following key insight: text optimization on an input image has shown promise in enabling smooth changes of an object toward the desired editing concept. However, this hinges on the condition that the model faithfully reproduces the input image and ensures that small changes in text embedding do not significantly alter the output image’s resemblance to the input image Kawar et al. (2023).
@@
Therefore, Imagic Kawar et al. (2023) additionally introduced model finetuning, ensuring the model specialized in reproducing the input image. Nevertheless, as this method causes overfitting and color distortion problems, we make a pivotal change by replacing the model fine-tuning with latent inversion. This change maintains the integrity of the original data distribution, preserving the editability of text optimization. Additionally, we exploit the reconstruction power of latent inversion further by inserting source text embedding exclusively during the initial sampling steps. Then, we transition to target text embedding for the subsequent sampling steps, namely timestep-aware text injection. This concept is inspired by the recent study that reveals the pivotal role of the initial sampling steps in determining the coarse structure and outlines of an image, highly correlated with the identity of input. By selectively injecting source and target prompts considering the role of sampling timesteps, we can explore the smooth transition between identity preservation and editability. This approach allows us to find a sweet spot between these two factors.
@@
Our simple yet effective approach effectively harmonizes with text optimization, achieving complex non-rigid editing without losing the original identity. It does not require additional inputs other than one input image and text prompts. Unlike Imagic, our training-free method offers the advantage of reusing a pre-trained diffusion model regardless of images. Also, our method generates more natural and aesthetically pleasing images than attention-based methods. We compare our edited results both quantitatively and qualitatively with other editing methods, showing superior editing performance on the Stable Diffusion in terms of identity preservation, editability, and aesthetic quality. Moreover, we prove each component of our method is effective through ablation experiments and show the applicability of our method on the different image domain such as Anything-v4.
@@ 2.1 Text-guided Image Manipulation
Text-guided image manipulation involves editing an image based on a given text prompt while preserving the original input image identity. In the realm of this technique, the predominant approach for preserving the original image identity involves the use of an attention mechanism. This mechanism integrates self or cross-attention maps from the input image into the generation process of an edited image Hertz et al. (2022); Tumanyan et al. (2023); Cao et al. (2023); Parmar et al. (2023). Additionally, other methods blend noise maps of the source image and target image, thereby maintaining the background from the source image Couairon et al. (2022); Avrahami et al. (2022); Huang et al. (2023b). Other directions include model finetuning Ruiz et al. (2023); Valevski et al. (2023); Kawar et al. (2023); Li et al. (2023) or text optimization Gal et al. (2022); Han et al. (2023); Kawar et al. (2023), which focus on adapting a model or target text to closely match the input image, even to the extent of overfitting. Furthermore, score distillation has been introduced for image editing, where the input image identity is continuously distilled into the target image during the editing process Hertz et al. (2023). Finally, several approaches Brack et al. (2023a); Tsaban and Passos (2023); Brack et al. (2023b) utilize semantic guidance by adding an edit term in the original classifier-free guidance formula Ho and Salimans (2022).
Figure 2: Ablation. We illustrate the impact of the following components in non-rigid editing: text optimization (Imagic/Ours), model fine-tuning (Imagic), inversion (Ours), and timestep-aware sampling (Ours). The leftmost image is an original input image. The rest of the images are edited according to the target text: “A photo of a white horse jumping” and “A photo of a couple hugging on a beach”. Text optimization minimally affects identity preservation. Model fine-tuning from Imagic enhances image resemblance to the input, often at the expense of color fidelity. Conversely, our latent inversion stabilizes the overfit tradeoff and avoids color distortion. To further improve identity preservation, we incorporate source text embedding during the initial sampling timesteps.
@@ 2.2 Diffusion Inversion
For real image editing, the concept of inversion has been widely used to uncover the latent representation that faithfully reconstructs the input real image. Inversion methods in the diffusion model have been evolved in order to improve both reconstruction and editing ability. DDIM inversion Song et al. (2020) excels at reconstructing the original image for unconditional image generation scenarios. However, it faces challenges in conditional cases where the use of classifier-free guidance (CFG) Ho and Salimans (2022) magnifies the errors, resulting in decreased editability and reconstruction quality. This drawback of DDIM gave rise to active research in inversion specialized for conditional image generation. Null-text inversion Mokady et al. (2023) improves the inversion quality by optimizing null text embedding $\emptyset$ at each diffusion timestep. On the other hand, Direct Inversion Ju et al. (2023) takes a different approach by disentangling source and target diffusion branches, eliminating optimization. This strategy boosts content preservation quality and edit fidelity within a short time. EDICT Wallace et al. (2023) offers a mathematically exact inversion solution through an alternative inversion process that involves a pair of coupled noise vectors. Edit-friendly inversion Huberman-Spiegelglas et al. (2023) strongly imprints the input image onto latent noise maps by making them correlated across timesteps, enabling fast and effective editing without the need for optimization. More explanations on the diffusion process, classifier-free guidance and inversion methods that we utilized are in Appendix.
@@ 3 Preliminary
Stable Diffusion Model. Diffusion models aim to estimate a data distribution via a latent variable model. They perform a forward process that gradually adds a Gaussian noise on a clean image at each timestep $t\in[1,...,T]$ until it reaches a noise distribution. Conversely, the reverse process inversely denoises the noise distribution in order to generate a final clean image. The Stable Diffusion model Rombach et al. (2022), an open-source model, conducts these processes on the latent space. Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, it encodes $\bm{x}$ into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ using a VQ-GAN Esser et al. (2021) encoder $\mathcal{E}$. The model is trained on this latent space, applying the forward process to $\bm{z}_{0}$. Likewise, $\bm{z}_{0}$ is obtained from the latent variable $\bm{z}_{T}$ using the reverse process. Finally, the generated image is computed by decoding $\bm{z}_{0}$ back to the image space ($\bm{x}=\mathcal{D}(\bm{z}_{0})$). Also, an important function of Stable Diffusion is its capability to facilitate text-guided image generation. This is achieved by inputting a text embedding $\tau_{\phi}(\bm{y})$, which is a text prompt $\bm{y}$ processed by CLIP Radford et al. (2021) text encoder $\tau_{\phi}$. The text-guided conditional model is trained using the following objective:
@@
where $f_{\theta}$ is a neural network with parameter $\theta$, $t\sim\texttt{Uniform}[1,T]$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\bm{z}_{t}$ is a noise added version of $\bm{z}_{0}$ using the forward process. Our goal is to improve non-rigid editing with Stable Diffusion, aiming to enhance the accessibility of a publicly available model. Detailed explanations of the forward and reverse process of diffusion models are available in the Appendix.
Challenges in Non-rigid Editing. Compared to rigid editing, such as global style transfer, image translation, and appearance transformation, non-rigid editing stands out as highly challenging. Specifically, rigid editing handles relatively simple color adjustments while maintaining the image structure intact, which model filters can easily capture. Conversely, non-rigid editing involves understanding and partially modifying an underlying structure of the original object while maintaining its identity. For that, the model should faithfully reconstruct the original input while selectively editing the structural components that align with the target prompt. Since faithful reconstruction has a trade-off relationship with diverse editing in most text-guided image generation models, non-rigid editing often lacks identity preservation. Therefore, improving the identity preservation ability of the model is the main issue in enhancing the non-rigid editing performance.
Figure 3: Overview. Our method sequentially performs (A) text optimization, (B) latent inversion, and (C) timestep-aware injection sampling. (A) The target text is encoded into an optimized embedding to match the input image. (B) Null-text inversion is employed for additional identity preservation. (C) The source text embedding $\bm{e}_{src}$ instead of an interpolated embedding $\bm{e}_{int}$ is injected during the initial sampling steps to anchor the image’s global structure more effectively.
Imagic for Non-rigid Editing. Imagic Kawar et al. (2023), a pioneering method for non-rigid editing, introduced a two-step process: text optimization and model fine-tuning. Firstly, they optimize the target text embedding to best match the input image. Specifically, the target text is encoded to a text embedding $\bm{e}_{tgt}\in\mathbb{R}^{N\times d}$ via a text encoder Raffel et al. (2020), where $N$ is the number of tokens in the target text and $d$ is the token embedding dimension. Following this, they freeze parameters of the pre-trained diffusion model $f_{\theta}$ and optimize the target text embedding $\bm{e}_{tgt}$ using the diffusion model objective:
@@
Finally, during the sampling phase, Imagic linearly interpolates the original target text embedding with the optimized one using Equation 3 and pass this through the model to generate an edited image. Here, it can control the degree of editing by varying an interpolation ratio $\alpha$. This aims to establish a smooth transition of text embedding toward the desired editing concept while maintaining the original identity.
@@
It is worth noting that Imagic is particularly effective when applied to Imagen Saharia et al. (2022), a large-scale generative model only accessible for business. It successfully applies the editing concept while faithfully maintaining the original identity and background of the original image. However, the same recipe encounters limitations when applied to Stable Diffusion, as mentioned in previous works Han et al. (2023); Brack et al. (2023b). Our experiments also affirm the unsatisfactory editing performance of Imagic on Stable Diffusion (Sec 5), with a noticeable deficiency in identity preservation.

### v3

# Daejeon, Republic of Korea BMVC Author Guidelines

@@ Abstract
Text-guided non-rigid editing involves complex edits for input images, such as changing motion or compositions within their surroundings. Since it requires manipulating the input structure, existing methods often struggle with preserving object identity and background, particularly when combined with Stable Diffusion. In this work, we propose a training-free approach for non-rigid editing with Stable Diffusion, aimed at improving the identity preservation quality without compromising editability. Our approach comprises three stages: text optimization, latent inversion, and timestep-aware text injection sampling. Inspired by the success of Imagic, we employ their text optimization for smooth editing. Then, we introduce latent inversion to preserve the input image’s identity without additional model fine-tuning. To fully utilize the input reconstruction ability of latent inversion, we suggest timestep-aware text injection sampling. This effectively retains the structure of the input image by injecting the source text prompt in early sampling steps and then transitioning to the target prompt in subsequent sampling steps. This strategic approach seamlessly harmonizes with text optimization, facilitating complex non-rigid edits to the input without losing the original identity. We demonstrate the effectiveness of our method in terms of identity preservation, editability, and aesthetic quality through extensive experiments.
@@ 1 Introduction
Among various image editing tasks, non-rigid editing involves changing the posture or composition of an object while preserving the background and object identity of the original image. It is useful for various applications, such as drawing comic books or videos, where gradual transitions in object actions or compositions over the input image are essential. This task is particularly more challenging than rigid editing, such as global style transfer and image translation, which focus on changing color or textures while retaining the image structure.
@@
The challenges of non-rigid editing become more severe when combined with Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], primarily related to preserving image identity. Prior works in image editing address this problem through various approaches, including injecting attention maps of the source image into the target image generation process [Hertz et al.(2022)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-or, Tumanyan et al.(2023)Tumanyan, Geyer, Bagon, and Dekel, Xiao et al.(2023)Xiao, Yin, Freeman, Durand, and Han, Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng, Wang et al.(2024)Wang, Liu, and Xu, Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi], model finetuning [Ruiz et al.(2023)Ruiz, Li, Jampani, Pritch, Rubinstein, and Aberman, Valevski et al.(2023)Valevski, Kalman, Molad, Segalis, Matias, and Leviathan, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], and target text optimization [Gal et al.(2022)Gal, Alaluf, Atzmon, Patashnik, Bermano, Chechik, and Cohen-or, Han et al.(2023)Han, Yang, Kwon, and Ye, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani]. However, the attention mechanism often leads to unnatural images as it forcefully combines the structure of the source image with the target, resulting in distorted objects and compositions. Also, model finetuning on an input image can lead to a severe overfitting problem and color distortion, especially when applied to Stable Diffusion.
@@
To this end, our work explores improving identity preservation for non-rigid editing with Stable Diffusion by incorporating a text optimization framework. We focus upon the following key insight: text optimization on an input image has shown promise in enabling smooth changes of an object toward the desired editing concept. However, this hinges on the condition that the model faithfully reproduces the input image and ensures that small changes in text embedding do not significantly alter the output image’s resemblance to the input image [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani].
@@
Therefore, Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani] additionally introduced model finetuning, ensuring the model specialized in reproducing the input image. Nevertheless, as this method causes overfitting and color distortion problems, we make a pivotal change by replacing the model fine-tuning with latent inversion. This change maintains the integrity of the original data distribution, preserving the editability of text optimization. Additionally, we exploit the reconstruction power of latent inversion further by inserting source text embedding exclusively during the initial sampling steps. Then, we transition to target text embedding for the subsequent sampling steps, namely timestep-aware text injection. This concept is inspired by the recent study that reveals the pivotal role of the initial sampling steps in determining the coarse structure and outlines of an image. By selectively injecting source and target prompts considering the role of sampling timesteps, we can explore the smooth transition between identity preservation and editability. This approach allows us to find a sweet spot between these two factors.
@@
Our simple yet effective approach effectively harmonizes with text optimization, achieving complex non-rigid editing without losing the original identity. Unlike Imagic, our training-free method offers the advantage of reusing a pre-trained diffusion model regardless of images. Also, our method generates more natural and aesthetically pleasing images than attention-based methods. We compare our edited results both quantitatively and qualitatively with other editing methods, showing superior editing performance on the Stable Diffusion in terms of identity preservation, editability, and aesthetic quality. Moreover, we prove each component of our method is effective through ablation experiments and show the applicability of our method on the different image domain such as Anything-v4.
@@ 2.1 Text-guided Image Manipulation
Text-guided image manipulation involves editing an image based on a given text prompt while preserving the original input image identity. In the realm of this technique, the predominant approach for preserving the original image identity involves the use of an attention mechanism. This mechanism integrates self or cross-attention maps from the input image into the generation process of an edited image [Hertz et al.(2022)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-or, Tumanyan et al.(2023)Tumanyan, Geyer, Bagon, and Dekel, Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng, Parmar et al.(2023)Parmar, Kumar Singh, Zhang, Li, Lu, and Zhu, Wang et al.(2024)Wang, Liu, and Xu, Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi]. Additionally, other methods blend noise maps of the source image and target image, thereby maintaining the background from the source image [Couairon et al.(2022)Couairon, Verbeek, Schwenk, and Cord, Avrahami et al.(2022)Avrahami, Lischinski, and Fried, Huang et al.(2023)Huang, Tang, Dong, Lee, and Xu]. Other directions include model finetuning [Ruiz et al.(2023)Ruiz, Li, Jampani, Pritch, Rubinstein, and Aberman, Valevski et al.(2023)Valevski, Kalman, Molad, Segalis, Matias, and Leviathan, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani, Li et al.(2023)Li, Huang, Ding, and Li] or text optimization [Gal et al.(2022)Gal, Alaluf, Atzmon, Patashnik, Bermano, Chechik, and Cohen-or, Han et al.(2023)Han, Yang, Kwon, and Ye, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], which focus on adapting a model or target text to closely match the input image, even to the extent of overfitting. Furthermore, score distillation has been introduced for image editing, where the input image identity is continuously distilled into the target image during the editing process [Hertz et al.(2023)Hertz, Aberman, and Cohen-Or]. Finally, several approaches [Brack et al.(2024)Brack, Friedrich, Hintersdorf, Struppek, Schramowski, and Kersting, Tsaban and Passos(2023), Brack et al.(2023)Brack, Friedrich, Kornmeier, Tsaban, Schramowski, Kersting, and Passos] utilize semantic guidance by adding an edit term in the original classifier-free guidance formula [Ho and Salimans(2021)].
@@ 2.2 Diffusion Inversion
For real image editing, the concept of inversion has been widely used to uncover the latent representation that faithfully reconstructs the input real image. DDIM inversion [Song et al.(2020)Song, Meng, and Ermon] excels at reconstructing the original image for unconditional image generation scenarios. However, it faces challenges in conditional cases where the use of classifier-free guidance (CFG) [Ho and Salimans(2021)] magnifies the errors, resulting in decreased editability and reconstruction quality. This drawback of DDIM gave rise to active research in inversion specialized for conditional image generation. Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or] improves the inversion quality by optimizing null text embedding $\emptyset$ at each diffusion timestep. On the other hand, Direct Inversion [Ju et al.(2023)Ju, Zeng, Bian, Liu, and Xu] takes a different approach by disentangling source and target diffusion branches, eliminating optimization. This strategy boosts content preservation quality and edit fidelity within a short time. EDICT [Wallace et al.(2023)Wallace, Gokul, and Naik] offers a mathematically exact inversion solution through an alternative inversion process that involves a pair of coupled noise vectors. Edit-Friendly inversion [Huberman-Spiegelglas et al.(2023)Huberman-Spiegelglas, Kulikov, and Michaeli] strongly imprints the input image onto latent noise maps by making them correlated across timesteps, enabling fast and effective editing without the need for optimization.
@@ 3 Preliminary
Stable Diffusion Model. Diffusion models aim to estimate a data distribution via a latent variable model. They perform a forward process that gradually adds a Gaussian noise on a clean image at each timestep $t\in[1,...,T]$ until it reaches a noise distribution. Conversely, the reverse process inversely denoises the noise distribution in order to generate a final clean image. The Stable Diffusion model [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], an open-source model, conducts these processes on the latent space. Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, it encodes $\bm{x}$ into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ using a VQ-GAN [Esser et al.(2021)Esser, Rombach, and Ommer] encoder $\mathcal{E}$. The model is trained on this latent space, applying the forward process to $\bm{z}_{0}$. Likewise, $\bm{z}_{0}$ is obtained from the latent variable $\bm{z}_{T}$ using the reverse process. Finally, the generated image is computed by decoding $\bm{z}_{0}$ back to the image space ($\bm{x}=\mathcal{D}(\bm{z}_{0})$). Also, an important function of Stable Diffusion is its capability to facilitate text-guided image generation. This is achieved by inputting a text embedding $\tau_{\phi}(\bm{y})$, which is a text prompt $\bm{y}$ processed by CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] text encoder $\tau_{\phi}$. The text-guided conditional model is trained using the following objective:
@@
where $f_{\theta}$ is a neural network with parameter $\theta$, $t\sim\texttt{Uniform}[1,T]$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\bm{z}_{t}$ is a noise added version of $\bm{z}_{0}$ using the forward process. Our goal is to improve non-rigid editing with Stable Diffusion, aiming to enhance the accessibility of a publicly available model.
Figure 1: Ablation on Imagic and Our method. The images are edited according to the target text: “A photo of a white horse jumping” and “A photo of a couple hugging on a beach”.
Imagic for Non-rigid Editing. Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], a pioneering method for non-rigid editing, introduced a two-step process: text optimization and model fine-tuning. Firstly, they optimize the target text embedding to best match the input image. Specifically, the target text is encoded to a text embedding $\bm{e}_{tgt}\in\mathbb{R}^{N\times d}$ via a text encoder [Raffel et al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu], where $N$ is the number of tokens in the target text and $d$ is the token embedding dimension. Following this, they freeze parameters of the pre-trained diffusion model $f_{\theta}$ and optimize the target text embedding $\bm{e}_{tgt}$ using the diffusion model objective:
@@
Finally, during the sampling phase, Imagic linearly interpolates the original target text embedding with the optimized one using Equation 3 and passes this through the model to generate an edited image. Here, it can control the degree of editing by varying an interpolation ratio $\alpha$. This aims to establish a smooth transition of text embedding toward the desired editing concept while maintaining the original identity.
@@
It is worth noting that Imagic is particularly effective when applied to Imagen [Saharia et al.(2022)Saharia, Chan, Saxena, Li, Whang, Denton, Ghasemipour, Gontijo Lopes, Karagol Ayan, Salimans, et al.], a large-scale generative model only accessible for business. It successfully applies the editing concept while faithfully maintaining the original identity and background of the original image. However, the same recipe encounters limitations when applied to Stable Diffusion, as mentioned in previous works [Han et al.(2023)Han, Yang, Kwon, and Ye, Brack et al.(2023)Brack, Friedrich, Kornmeier, Tsaban, Schramowski, Kersting, and Passos]. Our experiments also affirm the unsatisfactory editing performance of Imagic on Stable Diffusion (Sec 5), with a noticeable deficiency in identity preservation.
Figure 2: Overview of our method.